How to find the equation of the line that passes through the point (3,-2), and is parallel to the line with equation 3x-2y=7.
do I first find the slope? Then do I use the slope intercept form to get the equation?

Answer by richwmiller(17219)

Yes and yes.
slope is -a/b=-3/-2=3/2
-2=3/2*3+b
-2=9/2+b
-2-9/2=b
-4/2-9/2=b
-13/2=b
y=3/2*x-13/2
2y=3x-13
3x-2y=13
and the old equation was
3x-2y=7
check
2*-2=3*3-13
-4=9-13
-4=-4
ok